x(2x-1)-10=-2x(1-x)
Answer by stanbon(75887)

x(2x-1)-10=-2x(1-x)
2x^2-x-10 = -2x+2x^2
-x-10 = -2x
x=10
=========================
Cheers,
Stan H.